Lindiwe put 22 boxes in a storage unit each box weighed 5kg how can she use distributive propertyto write an expression to find the total weigh in the storage unit
Answer by fractalier(6550)

22*5 = (20+2)(5) = 20*5 + 2*5 = 100 + 10 = 110 kg